(a) Carbonate buffers are important in regulating the pH of blood at 7.40. what is the concentration ratio of CO2 (usually written H2CO3) to HCO3-in blood at pH = 7.40? H2CO3(aq)-HCO3-(aq) + H+(aq) Ka = 4.3 × 10-7 [H2CO3] / [HCO3-]- (b) Phosphate buffers are important in regulating the pH of intracellular fluids at pH values generally between 7.1 and 7.2. What is the concentration ratio of H2Po4 to HPO2 in intracellular fluid at pH = 7.15? H2PO4(aq)HPO42(aq)H(aq) Ka = 6.2 × 10-8
Answer #1

Answer:

a) H2CO3 (aq) \rightleftharpoons HCO3-(aq) + H+ (aq)

   Acid                  Conjugate base

pH of the buffer is given by Henderson-Hasselbalch equation as,

pH = pKa + log([Conjugate base]/[Acid])

For H2CO3 and HCO3- buffer,

pH = pKa (H2CO3) + log([HCO3-]/[H2CO3]) ..... (1)

Given data: pH of blood = 7.40,

Ka = 4.3*10-7. \Rightarrow pKa = -logKa = -log(4.3*10-7) = 6.37

Placing these values in eq.(1)

7.40 = 6.37 + log([HCO3-]/[H2CO3])

log([HCO3-]/[H2CO3]) = 7.40-6.37

log([HCO3-]/[H2CO3]) = 1.03

[HCO3-]/[H2CO3] = 101.03.

[HCO3-]/[H2CO3] = 10.7

On reciprocal,

[H2CO3] / [HCO3-] = 1 / 10.7

[H2CO3] / [HCO3-] = 0.09

Is the required ratio.

=========================================

b) H2PO4- (aq) \rightleftharpoons HPO42- (aq) + H+ (aq)

   Acid                  Conjugate base

pH of the buffer is given by Henderson-Hasselbalch equation as,

pH = pKa + log([Conjugate base]/[Acid])

For H2PO4- (aq) and HPO42- (aq)

pH = pKa (H2CO3) + log([HPO42-]/[H2PO4-]) ..... (1)

Given data: pH of blood = 7.15,

Ka = 6.2*10-8. \Rightarrow pKa = -logKa = -log(6.2*10-8) = 7.21

Placing these values in eq.(1)

7.15 = 7.21 + log([HPO42-]/[H2PO4-])

log([HPO42-]/[H2PO4-]) = 7.15-7.21

log([HPO42-]/[H2PO4-]) = -0.06

[HPO42-]/[H2PO4-] = 10-0.06.

[HPO42-]/[H2PO4-] = 0.87

On reciprocal,

[H2PO42-]/[HPO42-] = 1/0.87

[H2PO42-]/[HPO42-] = 1.15

Is the required ratio.

===================XXXXXXXXXXXXXXXXXXX====================